what is the x and y intercepts for the equation y=x-1
Answer by palanisamy(496)

The equation of the line is y=x-1
y-x = -1
To find the x intercept, put y=0
Then 0-x = -1
x=1
So the x intercept is (1,0)
To find the y intercept, put x=0
Then y-0 = -1
y= -1
So the x intercept is (0, -1)